(Part 1) An investor wants to have $1,000,000 when she retires in 20 years. If she...

  1. (Part 1) An investor wants to have $1,000,000 when she retires in 20 years. If she can earn 7% annual return on her investment the lump-sum she would need to invest today to reach her goal is closest to:
    1. $543,632
    2. $368.542
    3. $258,419
    4. $415.358
  1. (Part 2) An investment promises to pay $100 one year from today, $200 two years from today, and $300 three years from today. If the required rate of return is 14%, the value of the investment today is closest to:
    1. 444
    2. 333
    3. 222
    4. 555
Answer #1

Answer:

1]

PV = ? (amount which should be invested today)

FV = $ 1,000,000

t = 20 years and i = 7%

PV x (1 + 7%)20 = 1000000

PV = 1000000 / (1 + 7%)20

PV = 1000000 / 3.8696

PV = 258,419.002

Hence option C] $258,419 is correct option

2]

Present value (PV) ($) = 200 x P/F(14%, 1) + 300 x P/F(14%, 2) + 600 x P/F(14%, 3)

= 100 x 0.8772 + 200 x 0.7695 + 300 x 0.675

= 87.72 + 153.9 + 202.5

= 444.12

Hence option A]444 is correct option.